OpenStudy (anonymous):

2x to the second power - 1 divided by x to the second power + 3

OpenStudy (dumbcow):

\[\frac{2x^{2} - 1}{x^{2} +3}\]

OpenStudy (anonymous):

also when x = 1

OpenStudy (dumbcow):

??

OpenStudy (anonymous):

When x = 0, the answer is zero. Because, 1*1*1*1 = 1-1 = 0 If you have a zero for a numerator the value is zero.

OpenStudy (anonymous):

FYI, the fraction is 0/4.